A solution is prepared from 4.5743 g of magnesium chloride and 43.246 g of water. The vapor pressure of water above this solution is found to be 0.3630 atm at 348.0 K. The vapor pressure of pure water at this temperature is 0.3804 atm. Find the value of the van't Hoff factor i for magnesium chloride in this solution.
Answer #1

Answer:-

This question is solved by using the simple concept of relative lowering of vapor pressure using simple formula.

The answer is given in the image,

Ansuresi Massof Mydlo = 4.57439. Mass q coater = 43.24692 Por = 0.3630 po =0.3804 Molar mass of My Cl₂ = 24.3+ 2435.5 - = 95.22= n2 not m -0.0480 0.9026+0.0480 22= 0.0 1959 poops = 1 x po 0.3804-0.3630 = ixo.o1959 0.3804 0.04574 = ix0.01959 é = 2.33